Activity has been at an overall low level punctuated by a few short periods of pyroclastic flow generation from the dome above the Galway's Wall. Visibility has not been as good as the clear conditions of the past few days but some good views of the dome were obtained this morning.
Seismic activity has been at the same low level as yesterday. Two (2) volcano-tectonic earthquakes were recorded, and 27 rockfall signals. Two (2) long period events were also recorded. A period of continuous low-level rockfall activity from the dome above the Galway's Wall occurred between 7:30 am and 8:00 am. A burst of pyroclastic flow generation from the same area of the dome began at about 10:30 am and generated significant ash clouds which were blown to the WNW. None of these flow reached further than the Galway's Soufriere.
A series of helicopter flights this morning during the period of heightened pyroclastic flow generation allowed excellent views of the dome and the flows themselves. The area of recent dome growth has pushed further southward toward the Galway's Wall. The upper surface of this area is highly fractured and there are some high temperature gas emissions from some of the cracks. The chute through the Galway's Wall is almost completely filled with rock debris. Fumaroles in the area between the recent extrusion and the summit of the January 20 dome were emitting billowing steam clouds. The greyish colour of these clouds indicates that there is also some ash in them. Excellent views of the pyroclastic flows were possible. Run-out distances rarely reached further than the soufriere with most flows stopping in the area at the bottom of the Galway's Wall. The northern and eastern flanks of the dome complex remains unchanged but there are some precariously positioned blocks.
Processing of yesterdays GPS occupation of BIGNET has shown that there are no significant changes to any of the points making p the network. No EDM or COSPEC measurements were made today. The COSPEC runs made yesterday did not produce helpful results. This is attributed to the broad spread of the plume away from paths taken by the COSPEC instrument.
Despite the decrease in alert level, the volcano still remains dangerous. The dome is large, and a major collapse could start at any time. The activity today further shows that the activity can escalate with very little precursory activity. Daytime visits to zone C should be kept short, and nobody should enter zone A, which includes St Patrick's and Long Ground. People should remain alert and continue to use ash masks when necessary.
Professor Steve Sparks left Montserrat yesterday, after spending 5 weeks as the Chief Scientist of MVO. He has gone to the Seismic Research Unit, Trinidad to attend a meeting of Chief Scientists. Dr Simon Young is also attending the meeting, and will then travel to Montserrat to take up the Chief Scientist post. In the meantime, Dr Paul Jackson from the Seismic Research Unit is acting Chief Scientist.